OpenStudy (anonymous):

A gas of unknown molecular mass was allowed to effuse through a small opening under constant pressure conditions. It required 72 seconds (s) for 1.0 L of the gas to effuse. Under the identical experimental conditions, it required 28 s for 1.0 L of O2 gas to effuse. Calculate the molar mass of the unknown gas.

OpenStudy (anonymous):

\[\frac{ rate 1 }{ rate O_{2} } = \sqrt{\frac{ 32g }{ M1 }}\] rate 1 = 1L/72s = 0.01388 rate O2 = 1L/28s =0.0357\[\frac{32g}{(0.38864)^{2}} = M1 = 211.86 g\]